In my textbook these are called Perimeter Problems. I'll greatly appreciate any help I can get! If I can get help before Thurday November 4, it will be even more appriciated seeing as though I have a quiz on this chapter that evening!
One side of a triangle is three times the shortest side.
The third side is 7 feet more than the shortest side.
The perimeter is 62 feet. Find all three sides.

Answer by stanbon(75887)

One side of a triangle is three times the shortest side.
The third side is 7 feet more than the shortest side.
The perimeter is 62 feet. Find all three sides.
----------------
Let the shortest side be "x".
Then 3rd side is "x+7".
2nd side is "3x".
--------------------------
Perimeter = shortest + 3rd + 2nd
62 = x + x+7 + 3x
5x+7 = 62
5x = 55
x = 11
-----------------
shortest = 11 ft.
3rd side = x+7 = 18 ft
2nd side = 3x = 33 ft.
================================
Cheers,
Stan H.